Please show work for the following question:

Seasons Construction is constructing an office building under contract for Cannon Company and uses the percentage-of-completion method. The contract calls for progress billings and payments of $1,550,000 each quarter. The total contract price is $18,600,000 and Seasons estimates total costs of $17,750,000. Seasons estimates that the building will take 3 years to complete, and commences construction on January 2, 2018. At December 31, 2018, Seasons estimated that it was 30% complete with the construction, based on costs incurred. At December 31, 2019, Seasons Construction estimates that it is 75% complete with the building; however, the estimate of total costs to be incurred has risen to $18,000,000 due to unanticipated price increases. What is reported in the balance sheet at December 31, 2019 for Seasons as the difference between the Construction in Process and the Billings on Construction in Process accounts, and is it a debit or a credit? Difference between the accounts Debit/Credit

a) $4,225,000 Credit

b) $1,550,000 Debit

c) $1,100,000 Debit

d) $1,550,000 Credit

The answer is B) 1550000 Debit but why?
Answer #1

answer

B ) $ 1550000 Debit

Construction in process account = 18600000 * 75 %

= 13950000

Construction billing account balance = 1550000 * 4 *2

=12400000

Difference between the accounts = 13950000 - 12400000

= 1550000 debit